An electric ceiling fan is rotating about a fixed axis with an initial angular velocity magnitude of 0.300 rev/s . The magnitude of the angular acceleration is 0.885 rev/s2 . Both the the angular velocity and angular accleration are directed counterclockwise. The electric ceiling fan blades form a circle of diameter 0.740 m .

A.) Calculate the magnitude at of the tangential acceleration of a point on the tip of the blade at time t = 0.196 s .

Express the acceleration numerically in meters per second squared.

B.) Calculate the magnitude ar of the radial (or centripetal) acceleration of the point at the end of the fan blade.
Answer #1

A) tangential acceleration = angular acceleration*radius

= 0.885*2pi*0.74

= 4.115 m/s^2

B) Radial acceleration = w^2r = (w1 +alpha*t)^2*r

= [0.300*2pi+0.885*2pi*0.196]^2*0.74

= 6.55 m/s